I am STUCK! Please help me.
Can you solve by the addition method?
11x-8y=27
3x+5y=-7
Thanks.

Answer by stanbon(75887)

Let's say you want to get rid of the "x's"
Multiply the 1st equation by "3" and the 2nd equation by "-11":
33x-24y=81
-33x-55y=+77
Now add the equations to get:
-79y=158
y=-2
Substitute that back into either of the origiinal equations and
solve for "x":
3x+5(-2)=-7
3x-10=-7
3x=3
x=1
Solution is x=1;y=-2
Check it out in the "other" original equation:
11(1)-8(-2)=27
11+16=27
27=27
Cheers,
stan H.